OpenStudy (anonymous):

any1 can explain the principles in solving laplace equation in rectangular coordinates? i know i need to solve for 2 2nd order ODE but im not sure how to conclude for whether they are trivial or nontrivial.

OpenStudy (anonymous):

i let u(x, y) = F(x) G(y) so Uxx = F'' G Uyy = F G" then the equation become F"G + FG" = 0 i separate the variables and equate them to a separation constant, k so the 2 2nd order ODE are F''-kF=0 , G"+kG = 0 how do i proceed with the initial conditions given?

OpenStudy (anonymous):

i need to get 3 cases when k=0, k<0 and k>0. but i dont know how to conclude after applying the initial conditions into the cases.

OpenStudy (experimentx):

do you have relevant links?

OpenStudy (turingtest):

I always forget how to do this, so here is a hopefully helpful link http://tutorial.math.lamar.edu/Classes/DE/LaplacesEqn.aspx

OpenStudy (anonymous):

let say i have the case where k=0 for F"=0 i got F=Ax+B, and i substitube initial conditions to obtain F be trivial (F=0). for G"=0 i got G=Cx+D G(0)=0 : D=0, G=Cx G(24)=24: C=1, G=x ?? is it possible to obtain nontrivial solution for G when F is trivial?

OpenStudy (experimentx):

hold on for a while ... I'll work out few other probs and try to work on my copy.

OpenStudy (anonymous):

oh kay. thx alot

OpenStudy (anonymous):

my lecturer practically skipped this part ==

OpenStudy (anonymous):

@edr1c The devil is in the not-details, in the boundary conditions. Each set of B.C. allows/disallows other solutions

OpenStudy (anonymous):

The coordinates is in some sense "our imagination" - the equation is not influenced by the coordinates we choose. It IS influenced by the symmetry of its BOUNDARY, and of its BOUNDARY CONDITIONS

OpenStudy (anonymous):

So why do we choose the coordinates ? - To fit the symmetry of the boundary of course !